A 2 x 4 factorial experiment is conducted to compare yields of 4 varieties of soybeans that are planted in rows either 15 inches or 30 inches apart. Two plots of ground are randomly assigned to each combination of soybean variety and row spacing. The yields of soybeans (in bushels per acre) are as follows:

Rows 1 2 3 4
15" 45 46 47 46
46 46 48 43
30" 35 41 42 39
32 39 38 41

The partially completed ANOVA table is as follows:

Source df SS MS F
Total 319.75
Variety 41.25 13.75 5.0
Row spacing 225 225 81.8
Variety x row spacing 31.5
Error 22 2.75

Calculate the mean squares and then the F value for the variety x row spacing interaction.

Group of answer choices

F = 5.29

F = 1.96

F = 3.818

F = 16.36
Answer #1

Interaction degrees of freedom = (4 - 1)*(2 - 1) = 3

So,

MS (Variety x row spacing) = 31.5/3 = 10.5

Hence,

F = 10.5/2.75 = 3.818

Option C is correct.